OpenStudy (anonymous):

Two Martians, Splott and Fizzle, have solved the equation 2x + 4 = –3x + 14. Examine the work of Splott and Fizzle. Identify any errors in the Martians’ calculations and explain, using complete sentences, what corrections they should make. Create your own function to teach the Martians about functions. Your function must contain at least two different operations. Using complete sentences, prove to Splott and Fizzle that your function is a legitimate function. Using your function, explain to the Martians how to solve for f(3). Show your work and explain each step using complete sent

OpenStudy (anonymous):

Using complete sentences, describe to the Martians how to find the inverse of your function. The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2.

OpenStudy (anonymous):

Splott Fizzle 2x + 4 = –3x + 14 2x + 4 = –3x + 14 2x + 4 + 3x = –3x + 14 + 3x 2x + 4 – 2x = –3x + 14 – 2x 5x + 4 = 14 4 = 14 – 5x 5x + 4 + 4 = 14 + 4 4 – 14 = 14 – 5x – 14 5x = 18 –12 = –5x 5 x divided by 5 is equal to 18 divided by 5 –12 + 5 = –5x + 5 x = 18 over 5 –7 = x

OpenStudy (the_fizicx99):

Yish this is going to take sometime lol xD have you tried answering #1 ?

OpenStudy (anonymous):

Honestly I don't know how to.. I am literally brain dead when it comes to math. Putting letters in it didn't help either. :/ I have to have this completed by Friday along with other assignments, or I will be dropped.

OpenStudy (the_fizicx99):

Flvs too xD yeah I understand >.> I procrastinate myself :/ For #1 says to create a function, so it should look something like f(x) = 2x + 4

OpenStudy (anonymous):

exactly like that or? 0.o

OpenStudy (the_fizicx99):

For #2 a function is basically, whatever input you give it will have a unique y output, so in other words whatever you input you're going to be given an output. To find f(3) plug in 3 for every x into the equation, f(x) = 2x + 4 --> f(3) = 2(3) + 4 f(3) = 6 + 4 = 10 f(3) = 10

OpenStudy (the_fizicx99):

function *

OpenStudy (anonymous):

is that it? o.o

OpenStudy (the_fizicx99):

Oh wait, it says to create your own function! Lol make is simple for you, try f(x) = 2x + 6 >.> it's technically different xD The inverse function is \(\ f(x)^{-1} \) So basically: f(x) is another fancy way of saying y y = 2x + 6 1. exchange 'y' for 'x', so it looks like: x = 2y + 6 2. subtract 6 from both sides, x - 6 = 2y 3. divide both sides by 2 to isolate y 4. \(\ \dfrac{x-6}{2} = \dfrac{2y}{2} \) And that's your inverse function, just put f(x) back. So, \(\ \dfrac{x-6}{2} = f(x) \)

OpenStudy (anonymous):

what about the first part of the question?

OpenStudy (the_fizicx99):

It says to examine their work, the layout provided looks messy i can't tell if it's a part. Can you fix it into 2 separate comments?

OpenStudy (anonymous):

Examine the work of Splott and Fizzle. Identify any errors in the Martians’ calculations and explain, using complete sentences, what corrections they should make. Create your own function to teach the Martians about functions. Your function must contain at least two different operations. Using complete sentences, prove to Splott and Fizzle that your function is a legitimate function. Using your function, explain to the Martians how to solve for f(3). Show your work and explain each step using complete sent

OpenStudy (anonymous):

Using complete sentences, describe to the Martians how to find the inverse of your function. The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2.

OpenStudy (the_fizicx99):

No i mean the work they did

OpenStudy (anonymous):

oohhh okay!!

OpenStudy (the_fizicx99):

Yup ;p

OpenStudy (the_fizicx99):

@Emine<3

OpenStudy (anonymous):

splott 2x+4=-3+14 2x + 4 + 3x = –3x + 14 + 3x 5x + 4 = 14 5x + 4 + 4 = 14 + 4 5x = 18 –12 = –5x 5 x divided by 5 is equal to 18 divided by 5 x = 18 over 5

OpenStudy (anonymous):

Fizzle 2x + 4 = –3x + 14 2x + 4 – 2x = –3x + 14 – 2x 4 = 14 – 5x 4 – 14 = 14 – 5x – 14 –12 + 5 = –5x + 5 –7 = x

OpenStudy (anonymous):

is that better?

OpenStudy (the_fizicx99):

Yesh :>

OpenStudy (anonymous):

okay! :)

OpenStudy (the_fizicx99):

Splott 2x + 4 = -3x + 14 Subtract 4 from both sides, 2x = -3x + 10 add 3x to both sides, 5x = 10 /5 /5 x = 2 <-- what she should have gotten He failed when he added 4 to both sides, he should have subtracted it so they would cancel out. Anything below that is considered incorrect. I don't know where he got 5x = 18 - 12 = -5x >.> Fizzle - He failed miserably, his calculations are off, 4 - 14 = -10 not -12, because he messed up there, anything he continues to do using that wrong calculation will be incorrect. No clue how he go this either o.0 –12 + 5 = –5x + 5 ... I think he tried to add 5x to both sides. 5x is being multiplied not added, therefore the inverse of multiplication is division, he should have divided both sides by -5 if that was what he intended.

OpenStudy (anonymous):

well then.. Im glad some of us are gifted.. lol I would have terribly BSed that.

OpenStudy (the_fizicx99):

xD

OpenStudy (the_fizicx99):

Sill need help on the last ones?

OpenStudy (anonymous):

yes, sorry!! >.<

OpenStudy (the_fizicx99):

It's k xD no worries :> I believe this is the last one "The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2." Do you know function compositions?

OpenStudy (anonymous):

uhmm no?

OpenStudy (anonymous):

f(x) = 2x + 6 The inverse function is f(x)−1 y y = 2x + 6 1. x = 2y + 6 2. x - 6 = 2y 3. y 4. x−62=2y2 x−62=f(x) would this work as the created function?

OpenStudy (the_fizicx99):

hmm, kk Well technically no, assigning any number to 'x' will not give you the same answer because when they're 2 separate functions. Lemme double check

OpenStudy (anonymous):

hmmm okay..

OpenStudy (the_fizicx99):

Yeah pretty sure lol, always remember that it's being raised to a power :P \(\ f(x)^{-1} \)

OpenStudy (anonymous):

so is that how i would submit it? o.0

OpenStudy (the_fizicx99):

f(x) = 2x + 6 is a function yes, the inverse would be, \(\ f(x)^{-1} \rightarrow y = 2x + 6 \rightarrow x = 2y + 6 \rightarrow x - 6 = 2y \rightarrow \dfrac{x-6}{2} = \dfrac{2y}{2}\) So \(\ \dfrac{x-6}{2} = y \) or \( f(x) = \dfrac{x-6}{2} \) Now you create a new function, h(x) = ?

OpenStudy (anonymous):

uhm?

OpenStudy (the_fizicx99):

Create a new function now :P

OpenStudy (anonymous):

how would i do that?

OpenStudy (the_fizicx99):

I just gave you your inverse function for f(x) = 2x + 6 cx Anything you would like -- preferably easy lol Try something like h(x) = 4x

OpenStudy (the_fizicx99):

um, h(x) = 2x + 8 ? agree?

OpenStudy (anonymous):

yes!

OpenStudy (the_fizicx99):

kk sec

OpenStudy (anonymous):

okay!

OpenStudy (the_fizicx99):

ok, so using 2x + 8 h(2) = 2(2) + 8 h(2) = 4 + 8 h(2) = 12 Now take that result and plug it in f(x), f(x) = 2x + 6 f(h(2)) = f(12) = 30 understand?

OpenStudy (anonymous):

is that how I would submit this? o.0 or?

OpenStudy (the_fizicx99):

Ok just put the basic xD sole f(h(x)) and h(f(x)) and just say that they're \(\ aren't \) the same sometimes

OpenStudy (the_fizicx99):

solve *

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

if im annoying or wasting your time you don't have to help. you've already done a lot! I just don't understand anything.

OpenStudy (the_fizicx99):

Nah it's fine :p let's just finish the last question ;D

OpenStudy (the_fizicx99):

...This is the last question, right o.0

OpenStudy (anonymous):

I understand what you are trying to tell me its just that I dont understand what itll look like to submit it. you know?

OpenStudy (anonymous):

and yes lol